OpenStudy (anonymous):

No idea how to solve this: 4(rootsign)x+1-x=4

OpenStudy (anonymous):

This? \[4\sqrt{x+1} -x = 4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[4\sqrt{x+1} -x = 4\] \[\implies 4\sqrt{x+1} = 4 + x\] \[\implies 16(x+1) = (4+x)^2\] Can you take it from there?

OpenStudy (anonymous):

yes , u square on both sides and u'll get 2 roots. one of them is called an extreneous root. this occurs because, the given equation is linear(highest power of x is 1) and when u square both sides , u create a root( highest power now is 2). note: to eliminate this exrteneous root, substitute the value back in the original equation to see if its valid. solution for this equation is: x=8.

OpenStudy (anonymous):

I don't think the other root is extraneous. x=0 is valid.

OpenStudy (anonymous):

I think I can do this. I'm ultimately solving for x so by doing the distribution equation posted by polpak, I should get the answer of 0 and/or 8?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

question: when I see a number before the sqaureroot, do I also square that? Like in this one, is that how you got 16?

OpenStudy (anonymous):

\[4\sqrt{x+1} -x = 4\] Is not a linear equation.

OpenStudy (anonymous):

yea i was wrong :/ sorry, 0 and 8 are the roots.

OpenStudy (anonymous):

Yes, I squared both sides. \[(4\sqrt{x+1})^2 = 4^2(\sqrt{x+1})^2 = 16(x+1)\]

OpenStudy (anonymous):

Or more correctly \[16|x+1|\]

OpenStudy (anonymous):

ok...lightbulb is atarting to come on :) I'm seeing it.

OpenStudy (anonymous):

thank you!